# Two-dimensional nonlinear Schrödinger equations

### v1

@@
The paper deals with nonlinear Schrödinger equations of the general form, depending on time and two spatial variables, the potential and dispersion of which are specified by one or two arbitrary functions. The equations under consideration naturally generalize a number of related nonlinear partial differential equations that occur in various areas of theoretical physics, including nonlinear optics, superconductivity, and plasma physics. Two- and one-dimensional reductions are described, which lead the studied nonlinear Schrödinger equations to simpler equations of lower dimension or ordinary differential equations (or systems of ordinary differential equations). Using methods of generalized and functional separation of variables, a number of new exact solutions of two-dimensional nonlinear Schrödinger equations of the general form are found, which are expressed in quadratures or elementary functions. To analyze the equations under consideration, both Cartesian and polar coordinate systems are utilized. Special attention is paid to finding solutions with radial symmetry. The exact solutions obtained in this work can be used as test problems intended to assess the accuracy of numerical and approximate analytical methods for solving complex nonlinear PDEs of mathematical physics.
Keywords: two-dimensional nonlinear Schrödinger equations, exact solutions, solutions in quadratures, solutions in elementary functions, two- and one-dimensional reductions, methods of generalized and functional separation of variables, principle of structural analogy of solutions, nonlinear optics
@@ 1.2 Terminology: what is meant by exact solutions of nonlinear PDEs
(c) Solutions that are expressed in terms of solutions of ordinary differential equations (ODEs) or systems of such equations.
Various combinations of the solutions described in paragraphs (a)–(c) are also allowed. In cases (a) and (b) the exact solution can be presented in explicit, implicit or parametric form.
@@
In this paper, various modifications of the methods of generalized and functional separation of variables [15, 50, 51] and the principle of structural analogy of solutions [30, 52, 53] were used to find exact solutions and reductions of two-dimensional spatial generalizations of nonlinear equations (2) and (3), which contain one or two arbitrary functions.
@@ 2 Nonlinear Schrödinger equations with two spatial variables
2.1 Nonlinear Schrödinger equations under consideration and their properties
In this paper we will consider two-dimensional nonlinear Schrödinger equations of the general form
@@ 2.1 Nonlinear Schrödinger equations under consideration and their properties
where $w=w(x,y,t)$ is the desired complex-valued function of real arguments, $t$ is the time variable, $x$ and $y$ are the Cartesian coordinates, $\Delta$ is the Laplace operator (i.e. $\Delta w=w_{xx}+w_{yy}$), and $i^{2}=-1$. Equations (4) and (5) are two-dimensional generalizations of equations (2) and (3). The real functions $f(|w|)$ and $g(|w|)$ included in these nonlinear PDEs and characterizing the dispersion and potential will be considered arbitrary.
A group classification of the equation (4) was carried out in [54, 55, 56]. Some exact solutions of this equation in two-dimensional, three-dimensional and $n$-dimensional cases were obtained in [15, 57, 55] (see also [58, 59], where related PDEs were considered).
Properties of equations (4) and (5). Let $u(x,y,t)$ be a solution of equation (5). Then the functions
@@
where $A$, $C_{1}$, $C_{2}$, $C_{3}$, and $\beta$ are arbitrary real constants, are also solutions of this equation (the plus or minus signs in $w_{1}$ are chosen arbitrarily). For a simpler equation (4), the following property is added to the properties of (6): if $w(x,y,t)$ is a solution to the equation (4), then the function
@@
2.2 Transformation of the nonlinear Schrödinger equation to a system of real PDEs
$1^{\circ}$. For subsequent analysis, it is convenient to represent the desired function in exponential form
@@ 2.2 Transformation of the nonlinear Schrödinger equation to a system of real PDEs
Differentiating (8), we obtain
@@
Now we substitute (9) into (5), and then divide by $e^{i\varphi}$. Further equating the real and imaginary parts of the obtained relation to zero, we arrive at the following system of two real partial differential equations:
@@ 3.1 Two-dimensional traveling wave type solutions with constant amplitude
3.2 Solutions of the type of two-dimensional stationary solitons with an amplitude depending on one coordinate
@@
where $A$, $B_{1}$, $B_{2}$, and $C$ are arbitrary constants, and the function $r=r(x)$ is described by the nonlinear autonomous ODE
@@
It is easy to check that equation (13) admits the first integral
@@
which after resolution with respect to the derivative $h^{\prime}_{x}$ is reduced to a first-order ODE with separable variables. Integrating this equation, we obtain the general solution of equation (13) in implicit form
@@
The general solution of ODE (13), represented in quadratures in the implicit form (15), is difficult to study and interpret due to the presence of two arbitrary functions in it $f(r)$ and $g(r)$, which are included in the integrands. Therefore, here we use a simple and convenient for analysis semi-inverse approach developed in [28], based on introducing another arbitrary auxiliary function instead of the function $g(r)$ and directly specification of exact solutions in implicit form. Namely, we will consider the function $f(r)$ to be arbitrary, and the solution $r=r(x)$ of equation (13), taking into account the relation $h=rf$, will be specified using an arbitrary function $h=h(x)$ in implicit form
@@
Let us demonstrate with two specific examples how the described approach works in practice.
@@
For $f>0$ and $f^{\prime}_{r}\geqslant 0$ it follows from (21) that $r\to 0$ as $x\to\pm\infty$, i.e. the obtained solution decreases exponentially rapidly for large $x$. Moreover, the function $r=r(x)$ achieves its only maximum value $r_{m}$ at the point x=q/kx=-q/k; this value is determined from the algebraic (transcendental) equation $k\sqrt{2/c}=r_{m}f(r_{m})$.
Example 2. In relations (16)–(17) we substitute another rapidly decreasing function
@@
From (17) and (22)–(23) we get
@@
The resulting solution has properties qualitatively similar to the properties of the solution described in Example 1.
Remark 2. Searching for solutions of system (10) of a more general form $r=r(x)$, $\varphi=At+\theta(y)$ also leads to a solution (12).
3.3 Solutions of the type of two-dimensional stationary solitons with an amplitude depending on two coordinates

### v2

# arXiv:2512.07382v2[nlin.SI] 02 Mar 2026

@@
footnotetext: This article was published in the Open Journal of Mathematical Sciences, 2026, Vol. 10, pp. 173–194; doi: 10.30538/oms2026.0279.
For the first time, a nonlinear Schrödinger equation of the general form is considered, depending on time and two spatial variables, the potential and dispersion of which are specified by two arbitrary functions. This equation naturally generalizes a number of simpler nonlinear partial differential equations encountered in various fields of theoretical physics, including nonlinear optics, superconductivity, and plasma physics. Two- and one-dimensional reductions are described, which reduce the studied nonlinear Schrödinger equation to simpler equations of lower dimension or ordinary differential equations (or systems of ordinary differential equations). In addition to the general Schrödinger equation with two arbitrary functions, related nonlinear partial differential equations are also examined, in which the dispersion function is specified arbitrarily while the potential function is expressed in terms of it. For all considered classes of nonlinear PDEs, using the methods of generalized and functional separation of variables, as well as the semi-inverse approach and the principle of structural analogy of solutions, many new exact solutions have been found, which are expressed in terms of elementary or special functions, or in the form of quadratures. Both Cartesian and polar coordinate systems are employed to analyze the equations under consideration. Special attention is paid to finding solutions with radial symmetry. It is shown that the nonlinear Schrödinger equation, in which the functions defining the potential and dispersion are linearly related (one of these functions can be chosen arbitrarily), can be reduced to a two-dimensional nonlinear PDE that admits exact linearization. The exact solutions obtained in this work can be used as test problems intended for verifying the adequacy and assessing the accuracy of numerical and approximate analytical methods for solving complex nonlinear PDEs of mathematical physics.
Keywords: two-dimensional nonlinear Schrödinger equations, exact solutions, solutions in quadratures, solutions in elementary and special functions, two- and one-dimensional reductions, methods of generalized and functional separation of variables, semi-inverse approach, principle of structural analogy of solutions, nonlinear optics
@@ 1.2 Terminology: what is meant by exact solutions of nonlinear PDEs
(c) Solutions that are expressed in terms of solutions of ordinary differential equations (ODEs) or finite systems of such equations.
Various combinations of the solutions described in paragraphs (a)–(c) are also allowed. In cases (a) and (b) the exact solution can be presented in explicit, implicit, or parametric form.
@@
In this paper, various modifications of the methods of generalized and functional separation of variables [15, 50, 51], semi-inverse approach, and the principle of structural analogy of solutions [30, 52, 53] were used to find exact solutions and reductions of two-dimensional spatial generalizations of nonlinear equations (2) and (3), which contain one or two arbitrary functions.
@@ 2 Nonlinear Schrödinger equations with two spatial variables
2.1 Nonlinear Schrödinger equations under consideration and their properties. Preliminary description of the semi-inverse approach
$1^{\circ}$. In this paper we will consider two-dimensional nonlinear Schrödinger equations of the general form
@@
where $w=w(x,y,t)$ is the desired complex-valued function of real arguments, $t$ is the time variable, $x$ and $y$ are the Cartesian coordinates, $\Delta$ is the Laplace operator (i.e. $\Delta w=w_{xx}+w_{yy}$), and $i^{2}=-1$. Equations (4) and (5) are two-dimensional generalizations of equations (2) and (3). The real functions $f(r)$ and $g(r)$ with $r=|w|$ included in these nonlinear PDEs and characterizing the dispersion and potential will be considered rather arbitrary (it is assumed that the function $f$ is twice continuously differentiable, and the function $g$ is continuous).
A group classification of equation (4) with an arbitrary potential $g$ was carried out in [54]. Some exact solutions of this equation with two, three and $n$ spatial variables were obtained in [15, 54].
It is important to note that the nonlinear Schrödinger equation with two spatial variables (5), which includes two arbitrary functions $f$ and $g$, has not been considered before. Therefore, all the exact solutions of this equation obtained in the following sections are new.
$2^{\circ}$. In addition to the general Schrödinger equation with two arbitrary functions (5), we will also consider equations of the form (5) in which the function $f(r)$ will be set arbitrarily, and the second function $g(r)$ will be set parametrically by two relations
$rf(r)$ $=h(\xi),$ (6)
$g(r)$ $=M(\xi,r,h,h^{\prime}_{\xi},h^{\prime\prime}_{\xi\xi}),$
where $M$ is some given function of its arguments (it is derived from an ODE obtained as a result of a reduction), $\xi$ is a parameter (which can be a spatial variable or some ansatz of independent variables), $h=h(\xi)$ is a twice continuously differentiable function, the form of which can vary depending on a priori ideas about the structure of the solution. By inserting a specific invertible function $h(\xi)$ into (6), after eliminating $\xi$ we obtain a representation of the potential function in the form $g(r)=Q(r,f(r))$. In this case, the function $f(r)$ remains arbitrary, and the function $g(r)$ is expressed through it (further, using specific examples, it will be shown that the invertibility of the function $h(\xi)$ included in (6) is not a necessary condition for such a representation).
By specifying in (6) various concrete functions $h=h(\xi)$, we will obtain various classes of nonlinear Schrödinger equations depending on one arbitrary function $f(r)$. In particular, substituting $h(\xi)=\xi$ in (6), we get $g(r)=M(rf(r),r,rf(r),1,0)$.
The semi-inverse approach described above, in which one function of the equation under consideration is specified arbitrarily, and the second is expressed through it using relations (6), will be discussed in more detail in Section 3.2, and specific examples of its application will be given in Sections 3.3, 3.4, and 4.3). It should be noted that the exact solutions to nonlinear Schrödinger PDEs of the form (5), obtained further in this article, are new.
$3^{\circ}$. Properties of equations (4) and (5). Let ${\color[rgb]{1,0,0}\textrm{$w$}}(x,y,t)$ be a solution of equation (5). Then the functions
@@
where $A$, $C_{1}$, $C_{2}$, $C_{3}$, and $\beta$ are arbitrary real constants, are also solutions of this equation (the plus or minus signs in $w_{1}$ are chosen arbitrarily). For the simpler equation (4), the following property is added to the properties of (7): if $w(x,y,t)$ is a solution to the equation (4), then the function
@@
Remark 1. Symmetries and exact solutions of the nonlinear Schrödinger equation (4) with three spatial variables with a potential of the special form $g(|w|)=a+b|w|^{2}+c|w|^{4}$ were considered in [56, 57, 55] (for exact solutions of related PDEs with some given specific functions $g$, see also [15, 54, 55, 16, 58, 59]).
2.2 Transformation of the nonlinear Schrödinger equation to a system of two real PDEs
For subsequent analysis, it is convenient to represent the desired complex-valued function in exponential form
@@ 2.2 Transformation of the nonlinear Schrödinger equation to a system of two real PDEs
Using relation (9), we have
$f(|w|)w=f(r)re^{i\varphi}=he^{i\varphi},\qquad\ h=rf(r),$ (10)
Differentiating (9) and (10), we successively obtain
@@
$[f(|w|)w]_{x}$ $=(he^{i\varphi})_{x}=(h_{x}+ih\varphi_{x})e^{i\varphi},\quad[f(|w|)w]_{y}=(h_{y}+ih\varphi_{y})e^{i\varphi},$
$[f(|w|)w]_{xx}$ $=[(h_{x}+ih\varphi_{x})e^{i\varphi}]_{x}=[h_{xx}-h\varphi_{x}^{2}+i(2h_{x}\varphi_{x}+h\varphi_{xx})]e^{i\varphi},$
$[f(|w|)w]_{yy}$ $=[(h_{y}+ih\varphi_{y})e^{i\varphi}]_{x}=[h_{yy}-h\varphi_{y}^{2}+i(2h_{y}\varphi_{y}+h\varphi_{yy})]e^{i\varphi},$
@@
Now we substitute (11) into (5), and then divide all terms by $e^{i\varphi}$. Further equating the real and imaginary parts of the obtained relation to zero, we arrive at the following system of two real partial differential equations:
@@ 3.1 Two-dimensional traveling wave type solutions with constant amplitude
3.2 Closed-form solution, periodic in time and in one spatial variable, with amplitude dependent on the other spatial variable
@@
where $A$, $B$, and $C$ are arbitrary constants, and the function $r=r(x)$ is described by the nonlinear autonomous ODE
@@
In what follows, it is assumed throughout that the potential function $f=f(r)$ satisfies the following conditions:
$f\geqslant 0,\qquad\ f^{\prime}_{r}\geqslant 0,\qquad\ f\not\equiv 0\qquad\ \text{for}\ \ r\geqslant 0.$ (16)
Let us now multiply all terms of equation (15) by $2h^{\prime}_{x}$, and then integrate with respect to $x$. Taking into account the relations $2h^{\prime}_{x}h^{\prime\prime}_{xx}=[(h^{\prime}_{x})^{2}]^{\prime}_{x}$ and $2hh^{\prime}_{x}=(h^{2})^{\prime}_{x}$, as a result we obtain
@@
where $C_{1}$ is an arbitrary constant. Since the function $h$ must be real, the right-hand side of the relation (17) must be nonnegative, which leads to the condition
$B^{2}h^{2}+2\int[Ar-rg(r)]dh+C_{1}\geqslant 0.$
The first integral (17), after resolution with respect to the derivative $h^{\prime}_{x}$, is reduced to a first-order ODE with separable variables. Integrating this equation, we obtain the general solution of equation (15) in implicit form
@@
The general closed-form solution of the ODE (15), represented in quadratures in the implicit form (18), is difficult to study and utilize due to the presence of two arbitrary functions $f(r)$ and $g(r)$, which are included in the integrands.
Note that for the defining functions of power-law form
$f(r)=\beta r^{\sigma},\qquad\ g(r)=\gamma r^{\sigma},$
where $\beta>0$, $\gamma>0$, and $\sigma>0$ are free parameters, solution (18) with $A>0$, $C_{1}=0$, and $\beta B^{2}>\gamma$ can be expressed explicitly in terms of elementary functions by the following formulas:
$r(x)$ $=\left[\frac{2A(\sigma+1)}{(\sigma+2)(\beta B^{2}-\gamma)}\sinh^{2}\left(\frac{\sigma\sqrt{\beta B^{2}-\gamma}}{2\sqrt{\beta}\,(\sigma+1)}(C_{2}\pm x)\right)\right]^{1/\sigma}$ $\text{if}\ \ \beta B^{2}>\gamma,$
$r(x)$ $=\left[\frac{2A(\sigma+1)}{(\sigma+2)(\gamma-\beta B^{2})}\sin^{2}\left(\frac{\sigma\sqrt{\gamma-\beta B^{2}}}{2\sqrt{\beta}\,(\sigma+1)}(C_{2}\pm x)\right)\right]^{1/\sigma}$ $\text{if}\ \ \beta B^{2}<\gamma,$
$r(x)$ $=\left[\frac{A\sigma^{2}}{2\beta(\sigma+1)(\sigma+2)}\,(C_{2}\pm x)^{2}\right]^{1/\sigma}$ $\text{if}\ \ \beta B^{2}=\gamma.$
Remark 3. Searching for solutions of system (12) of a more general form $r=r(x)$, $\varphi=At+\theta(y)$ also leads to solution (14).
3.3 Solutions, periodic in time and in one spatial variable, with amplitude dependent on the other spatial variable. Semi-inverse approach
$1^{\circ}$. Simplest version of the semi-inverse approach. Now we describe the simple and convenient for analysis semi-inverse approach developed in [28], which is based on introducing another arbitrary auxiliary function instead of the function $g(r)$ and directly specifying exact solutions in implicit form. Namely, we will consider the function $f(r)$ as arbitrary, and the solution $r=r(x)$ of equation (15), taking into account the relation $h=rf$, will be given in terms of an auxiliary function $h=h(x)$ in the implicit form
@@
Setting in (19)–(20) various specific functions $h=h(x)$, we will obtain different classes of nonlinear Schrödinger equations, depending on one arbitrary function $f(r)$, which admit exact solutions in the implicit form (19).
Note that in many cases it is not necessary to require the auxiliary function $h(x)$ to be monotonic (invertible), since it is not infrequently possible to directly express the second derivative of $h_{xx}^{\prime\prime}$ in terms of the function $h$ itself as
$h^{\prime\prime}_{xx}=Q(h).$ (21)
Substituting relation (21) into (20) and taking into account the connection $h=rf$, we obtain the following representation for the potential function:
$g(r)=A+B^{2}f(r)-r^{-1}Q(rf(r)),$ (22)
which determines the specific form of the nonlinear PDE (5).
Under conditions (16) and $h(x)\geqslant 0$, the implicit relation (19) defines a single-valued function $r=r(x)$.
Let us briefly summarize the above: in the semi-inverse approach, the solution $r=r(x)$ of the reduced ODE (15) is implicitly prescribed in the form (19), while the form of the original PDE (5) is determined by the potential function (22). It is important to note that when using this method, the dispersion function $f(r)$ in the equation and its solution remains arbitrary.
Let us demonstrate with four specific examples how the semi-inverse approach works in practice.
@@
If conditions (16) are satisfied, it follows from (26) that $r\to 0$ as $x\to\pm\infty$, i.e. the obtained solution decreases exponentially rapidly for large $x$. Moreover, the function $r=r(x)$ achieves its only maximum value $r_{m}$ at the point x=q/kx=-q/k; this value is determined from the algebraic (transcendental) equation $k\sqrt{2/c}=r_{m}f(r_{m})$.
For a power-law dispersion function
$f(r)=\beta r^{\sigma},$ (27)
where $\beta>0$ and $\sigma\geqslant 0$ are arbitrary parameters, equation (25) takes the form
$iw_{t}+\beta\Delta(|w|^{\sigma}w)+(a+b\beta|w|^{\sigma}+c\beta^{3}|w|^{3\sigma+2})w=0,$
and its exact solution is described by formulas (9) and (14), where the amplitude can be represented in explicit form
$r(x)=\biggl[\frac{k\sqrt{2/c}}{\beta\cosh(kx+q)}\biggr]^{\!\textstyle\frac{1}{\sigma+1}}.$
Example 2. In relations (19)–(20) we substitute another exponentially decreasing function
@@
From (20) and (28)–(29), taking into account (19), we obtain
@@
The resulting solution (31) has properties qualitatively similar to the properties of solution (26) described in Example 1.
Example 3. Let us now take a more rapidly decreasing function of the form
$h(x)=k\exp(-\lambda x^{2}),$ (32)
where $k$ and $\lambda$ are positive free parameters. Differentiating (32), we find
$h^{\prime\prime}_{xx}=-2k\lambda\exp(-\lambda x^{2})-4k\lambda^{2}x^{2}\exp(-\lambda x^{2})=-h(4\lambda\ln h+2\lambda-4\lambda\ln k).$ (33)
Substituting (32) and (33) into (20) and taking into account (19), we have
$g(r)=a+bf(r)+cf(r)\ln[rf(r)],\qquad\ a=A,\qquad b=B^{2}+2\lambda-4\lambda\ln k,\qquad c=4\lambda.$
$iw_{t}+\Delta[f(|w|)w]+[a+bf(|w|)+cf(|w|)\ln(|w|f(|w|))]w=0,$
where $f(r)$ is an arbitrary function, has an exact solution described by formulas (9) and (14), in which the function $r=r(x)$ is given by the implicit relation
$k\exp(-\lambda x^{2})=rf(r).$ (34)
Solution (34) has properties qualitatively similar to the properties of the solution described in Example 1.
For the dispersion function of power-law form (27), the amplitude from (34) can be expressed explicitly as
$r(x)=\biggl(\frac{k}{\beta}\biggr)^{\!\textstyle\frac{1}{\sigma+1}}\exp\biggl(-\frac{\lambda}{\sigma+1}x^{2}\biggr).$
Example 4. In relations (19)–(20) we substitute the periodic function
$h(x)=a+b\sin(kx+c),$ (35)
where $a$, $b$, $c$, and $k$ are free parameters ($a>b>0$). Taking into account that
$h^{\prime\prime}_{xx}=-bk^{2}\sin(kx+c)=ak^{2}-k^{2}h,$ (36)
and using (20) and (35)–(36) we find the potential
$g(r)=A-ak^{2}r^{-1}+(B^{2}+k^{2})f(r).$ (37)
Thus, we arrive at the nonlinear Schrödinger equation
$iu_{t}+[f(|u|)u]_{xx}+[A-ak^{2}|u|^{-1}+(B^{2}+k^{2})f(|u|)]u=0,$ (38)
which has an exact solution described by the formulas (9) and (14), where the amplitude $r=r(x)$ is given by the implicit relation
$a+b\sin(kx+c)=rf(r).$
$2^{\circ}$. More complex version of the approach based on differential constraints. This modification of the semi-inverse approach uses solutions defined from suitable auxiliary differential equations [29]. Namely, instead of an auxiliary function $h(x)$, we first introduce an auxiliary integrable ODE of the form (21), which is simpler than equation (15). In this approach, the function $f(r)$ is considered arbitrary, whereas the function $Q(r)$ is chosen so that the solution of the ODE (21) can be expressed in terms of elementary or special functions as $h=h(x)$. In this case, the potential function is expressed through $f(r)$ and determined by formula (22). The solution of the corresponding ODE (15) is described by the implicit relation (19). Note that in the described modification of the semi-inverse approach, the auxiliary ODE (21) can be interpreted as a differential constraint (see, for example, [15, 50, 53]).
Let us use a specific example to demonstrate how modified approach works in practice.
Example 5. Let us take a differential constraint with quadratic nonlinearity
$h^{\prime\prime}_{xx}=ah^{2},$ (39)
which is obtained by substituting the function $Q(h)=ah^{2}$ into (21). The general solution of equation (39) is expressed in terms of the Weierstrass function by the formula (see, for example, [60]):
$h(x)=\frac{6}{a}\wp(x+C_{1},0,C_{2}),$
where $C_{1}$ and $C_{2}$ are arbitrary constants. In this case, the potential function is defined by the expression
$g(r)=A+B^{2}f(r)-arf^{2}(r),$
where $f(r)$ is an arbitrary function specifying the dispersion. The exact solution of the ODE for the amplitude (15) admits a representation in implicit form
$\frac{6}{a}\wp(x+C_{1},0,C_{2})=rf(r).$
3.4 Time-periodic solutions with an amplitude depending on two spatial variables